For the circuit shown, the switch has been open for a long time. Explain about charges, currents, and voltages. Now, the switch closed at t1. Explain what is now true about the charges, currents, and voltages. Did the charge in C2 increase, or decrease?

For the circuit shown, the switch has been open fo
For the circuit shown, the switch has been open fo
Answer #1


after a long time Switch is closed ast S1,

Both Capactors will be at Sereis Combination and Both Resistors

so

Rnet = R1+R2 = 680+220 = 900 ohms

Cnet = C1C2/(C1/+C2)

Cnet = 220* 330/(220+330)

Cnet = 132 uF

so

intial charge Qi = CV = 132e -6 * 72    = 9.5 mC

Current i = V/R = 72/(900) = 0.08 Amps

initial Volatge   72 Volts

--------------------------------------


yes when switch is closed,

growth of charge takes place across the terminals of the Capacitor

so across C2 in increases